OpenStudy (anonymous):

help! Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = -14 x = -17

OpenStudy (anonymous):

is it (x+20)/2=3x or x+20/2=3x

OpenStudy (therealmeeeee):

i got this one

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

which one

OpenStudy (therealmeeeee):

x + 20/2 = 3x multiple both sides by 2 to get rid of the 2 2(x+20/2) = 3x(2) the 2s cancel each other out on the left side x+20=6x you want to get the x's on the same side so subtract x from each side 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

lol thanks @TheRealMeeeee

OpenStudy (therealmeeeee):

=4

OpenStudy (therealmeeeee):

yup anything for u

OpenStudy (anonymous):

lol :)

OpenStudy (therealmeeeee):

:)

OpenStudy (anonymous):

x+20/2=3x multiply both sides by 2 2(x+20/2)=(3x)2 2x+20=6x -2x -2x 20=4x divide both sides by 4 5=x

OpenStudy (anonymous):

@TheRealMeeeee you would be right if it was the other equation

OpenStudy (anonymous):

ookkkk

OpenStudy (anonymous):

wait how is it 5 its not a choice

OpenStudy (therealmeeeee):

i kniw i was right

OpenStudy (therealmeeeee):

u was wrong

OpenStudy (therealmeeeee):

@sodukogirl13 u was wrong......... was 4 a choice

OpenStudy (therealmeeeee):

@lulu22 was 4 a choice

OpenStudy (anonymous):

lol truuu

OpenStudy (anonymous):

yeah i put the choice up top

OpenStudy (therealmeeeee):

ohhhh

OpenStudy (anonymous):

no I wasn't but I guess if lulu wants to go with ur answer he/she may

OpenStudy (anonymous):

but 5 is not a choice and i am a girlll

OpenStudy (therealmeeeee):

shes a girll

OpenStudy (anonymous):

well I didn't know if u were a girl or a boy which is why I put he/she

OpenStudy (therealmeeeee):

a good lookin girl tbh